# Scaling Behavior Cloning Improves Causal Reasoning: An Open Model for Real-Time Video Game Playing

### v1

@@ Abstract
Behavior cloning is enjoying a resurgence in popularity as scaling both model and data sizes proves to provide a strong starting point for many tasks of interest. In this work, we introduce an open recipe for training a video game playing foundation model designed for inference in realtime on a consumer GPU. We release all data (8300+ hours of high quality human gameplay), training and inference code, and pretrained checkpoints under an open license. We show that our best model is capable of playing a variety of 3D video games at a level competitive with human play. We use this recipe to systematically examine the scaling laws of behavior cloning to understand how the model’s performance and causal reasoning varies with model and data scale. We first show in a simple toy problem that, for some types of causal reasoning, increasing both the amount of training data and the depth of the network results in the model learning a more causal policy. We then systematically study how causality varies with the number of parameters (and depth) and training steps in scaled models of up to 1.2 billion parameters, and we find similar scaling results to what we observe in the toy problem.
@@
11footnotetext: This paper is based, in part, on a prior paper Yue et al. (2025).22footnotetext: All data, code, model checkpoints and videos of game playing are available at the accompanying website https://elefant-ai.github.io/open-p2p/.
@@ 1 Introduction
In parallel, recent advances in large language models (LLMs), such as ChatGPT (Brown et al., 2020), have brought general-purpose AI into everyday use. Most commercial LLM systems have since evolved into multimodal visual language models (VLMs) that accept image inputs. Despite this progress, deploying large multimodal models for real-time control remains challenging on consumer-grade hardware. Even ignoring latency and cost constraints, current VLMs perform poorly at game control; for example, none are able to complete the first level of the 1996 shooter Quake (Zhang et al., 2025).
In this paper, we train a single model capable of playing a range of 3D video games using raw image observations and producing keyboard and mouse actions in real time on a consumer-grade GPU. Our approach leverages behavior cloning with a large-scale dataset collected across a diverse set of games. We release all training and inference code, as well as the dataset (an example is shown in Figure 1), under open licenses. We demonstrate that the model can play simple games that do not require a high level of planning with a high level of competence.
@@
Causal confusion occurs due to non-causal correlations in the data that can result in the policy learning to predict the action using these non-causal correlates. An example, taken from De Haan et al. (2019), is that the policy may learn to apply the brakes when it sees the brake lights, since these are highly correlated. Obviously, such a non-causal policy performs poorly.
We show, both in a simple toy example and in testing our model at different data and model sizes, that, in practice, in our setting, scaling both the model and dataset size results in a model that more reliably focuses on causal signals. This suggests that one practical solution to issues of causality in behavior cloning is simply scaling up both the model size and data size and diversity.
@@ 2.1 Annotated data
We collect a large-scale, high-quality dataset of human gameplay spanning a diverse set of popular 3D video games. The complete list of games is provided in Appendix A.1. Gameplay is recorded by experienced players who are instructed to capture only active gameplay segments (e.g., excluding lobby or waiting periods). Annotators use a variety of hardware, monitor sizes and resolutions, mouse sensitivities, and gameplay styles, which provide a diverse set of gameplay videos.
@@
After filtering for quality (see Appendix A.3 for details), the dataset comprises over 8,3008{,}300 hours of high-quality human gameplay, corresponding to more than $650$ million image–action pairs. The distribution of recording hours across games is shown in Appendix A.2
@@ 2.1.1 Text-Annotated Data
Text annotations were generated retrospectively using a commercial VLM. The VLM was prompted to review gameplay segments and infer plausible instructions that could have guided the human player at specific timestamps. While VLMs are not capable of playing games at a high level (Zhang et al., 2025), we found that they produce high-quality instructional descriptions when analyzing gameplay videos offline.
A key challenge in this process is that commercial VLMs typically operate on temporally downsampled video (often around 1 Hz), which is insufficient for text annotation in fast-paced games. To address this limitation, we deliberately prompt the VLM to generate temporally extended, goal-oriented instructions (e.g., “move toward the skull gate”) rather than instantaneous commands (e.g., “turn left”). We further design the prompt to suppress repetitive, high-frequency events, such as continuous shooting, by annotating such behaviors only once per contiguous segment.
@@
Figure 2: (2(a)) Architecture of P2P. The core policy transformer and action decoder are both decoder-only transformers. Each timestep begins with a text token $t_{i}$. Since many frames do not contain a text annotation there is a default text token $t_{null}$ used on these frames. This is followed by image token(s) from video frame $o_{i}$ followed by a learnable “reasoning” token $k_{i}$ that grants the model extra computation. The policy transformer then outputs a single action prediction token $a_{in}$. A smaller transformer, the action decoder, then auto-regressively transforms and samples the single action prediction token into the full action space. Then the true action tokens $a_{i}$ are input so that $a_{in}$ at time $i+1$ can attend to the true action tokens from time ii. (2(b)) Attention mask used in our transformer policy (green denotes 11 and gray 0). This custom mask ensures the action prediction token $a_{in}$ at time $i$ cannot attend to the ground truth action at time ii. Note that no other tokens attend to $a_{in}$ to stabilize the training process.
@@ 2.1.2 Correction Data
To collect correction trajectories, we deploy a trained policy to interact with the game environment while a human annotator monitors its behavior. When the policy encounters out-of-distribution situations (e.g., becoming stuck or exhibiting degenerate behavior), the annotator temporarily took control to guide the agent back to a valid state. Control is then returned to the policy. For behavior cloning we used only the human actions and masked out the loss for actions taken by the agent. These correction trajectories are mixed with the original annotated data during training and constitute less than 1% of the total annotated dataset.
@@ 2.1.3 Simple Benchmark Environments
In addition to commercial games, we collected data from two lightweight 3D environments we designed for automated evaluation: Hovercraft and Simple-FPS (see Appendix A.5). These environments are fully programmatic, allowing precise control over game state and difficulty, which provide a fair and controlled benchmark for comparing model performance (see Section 4).
@@ 3 Policy Model
We present a multimodal action policy model, which we refer to as Pixels2Play (P2P). P2P is a text-conditioned policy that takes visual observations and optional textual instructions as input, and outputs low-level keyboard and mouse actions. The model builds upon the transformer-based policy architecture introduced in Yue et al. (2025), but introduces substantial extensions to support text conditioning and to improve online performance by explicitly conditioning the backbone transformer on ground-truth action tokens during training.
A primary design constraint is that the model must operate in real time (20 Hz) on high-end consumer GPUs (e.g. NVIDIA RTX 5090), enabling deployment on end-user hardware. To satisfy this requirement, we trained a lightweight, decoder-only transformer backbone from scratch rather than fine-tuning a large pretrained VLM, as explored in prior work (Kim et al., 2024). This design offers two advantages: (i) it enables a custom image tokenization pipeline that produces a small number of image tokens, allowing the model to attend to longer temporal histories, which is important for games that require long-term memorization; and (ii) the reduced model size and custom architecture ensures fast inference and compatibility with compilation and optimization techniques.
We evaluated several image encoders, including MagVitV2 (Luo et al., 2024), DINOv2 (Oquab et al., 2023), IBQ (Shi et al., 2025), and CosmosTokenizer (Agarwal et al., 2025). We found that EfficientNet (Tan and Le, 2019), following Pearce and Zhu (2022), provides the best trade-off between representation quality and computational efficiency. Specifically, we use the first six layers of EfficientNet, followed by a linear projection into a small set of visual tokens (typically $N_{i}\in\{1,\dots,4\}$). Unfreezing the image encoder during training consistently improves performance compared to freezing it (Appendix C.1). The model operates directly on the resized raw pixel inputs (192 pixels $\times$192 pixels).
Much prior work focused on single-game agents and adopts reduced action spaces that can be modeled as a single categorical variable (Baker et al., 2022; Pearce and Zhu, 2022). In contrast, our setting requires a unified policy capable of operating across many games, necessitating a much larger action space comprising the full keyboard and mouse input domain. We allowed up to four simultaneous key presses and two concurrent mouse actions, making treating the action as a single one-hot combinatorial impractical.
@@
At each timestep, the policy transformer consumes image tokens ($o_{i}$), a text-conditioning token ($t_{i}$), ground-truth action tokens ($a_{i}$), and a single action prediction token ($a_{\text{in}}$). We further introduce an optional “thinking” token $k_{i}$, which provides the model with an additional reasoning step prior to action prediction. The resulting token count per timestep is $N_{i}+N_{a}+2$. Each token is augmented with a learned type embedding indicating its role (image, text, reasoning, ground-truth action, or action prediction). Rotary positional embeddings (Su et al., 2024) are applied at every transformer layer. During inference, we employ key–value caching with sliding-window attention to bound memory usage.
@@
Behavior cloning often suffers from causal confusion (De Haan et al., 2019). One particular case of this, which becomes worse at higher frequencies, is the model learns to copy previous actions instead of attending to visual inputs (Wen et al., 2020). Although this issue can be avoided by excluding ground-truth actions from the policy inputs, we find that conditioning the policy on ground-truth action tokens produces more human-like behavior than conditioning on visual observations alone. In particular, the policy learns to sustain actions over multiple frames before switching, closely resembling human gameplay dynamics. Qualitative differences in behavior are illustrated in the demonstration videos [^3] Under Action Conditioning section on the website..
@@ 3.0.1 Leveraging unlabeled data
Figure 3: (3(a)) Gap induced by video compression without data augmentation. We measured this gap by comparing model outputs on raw frames (inference) and resized frames (training). An irreducible gap arises from lossy video compression during data collection. The gap was smaller with RGB than YUV encoding and increases as compression quality degrades (lower file size). The x-axis qp denotes the quantization parameter, where larger values indicate lower quality. Data augmentation (3(b)) mitigates this gap when reasonable compression quality is used.
@@ 3.1 Mitigating the Training–Inference Gap
Early experiments revealed a substantial performance gap between offline evaluation and online deployment. We traced this gap to discrepancies between training and inference inputs arising from video recording, compression, and resizing. For practical latency and storage considerations, videos undergo two lossy processing steps during data collection: (1) compression and upload on the annotator side, and (2) resizing to $192\times 192$ for model processing. During inference resizing occurs but no compression takes place.
We observed a substantial divergence in model outputs when using uncompressed raw frames versus resized training frames, which leads to deceptively strong offline metrics but poor online performance. We measured this gap by collecting a small number of uncompressed videos. We then compressed the videos using differing compression options and compared the trained model probabilities on the uncompressed versus compressed output. Although compression at the annotation stage is unavoidable (which causes the irreducible gap), we find that the choice of color space during resizing plays a critical role: RGB encoding yields a smaller training–inference gap than YUV encoding (Figure 3(b)). Unfortunately, NVIDIA hardware encoders support only the YUV color space; therefore, we adopt a mixture of QP values to balance encoding speed and encoding quality.
@@
Data augmentation also substantially reduces the training–inference gap. We applied mild spatial transformations, color perturbations, Planckian jitter, ISO noise, random gaussian or motion blur, sharpening, and translation during training. As shown in Figure 3(b), these augmentations significantly reduced the discrepancy and improved online performance. Consequently, all experiments in this work employ data augmentation. We believe further improvements to reduce the training-inference gap via targeted data augmentation are an interesting area for future work.
Finally, discretization of the mouse action space can lead to overly aggressive or overly conservative mouse movements. Following Pearce and Zhu (2022), we discretized mouse actions using quantile-based bins, which provide fine resolution near zero but coarse resolution in the tails. To improve robustness, we fit truncated normal distributions to the x and y axes of the mouse action space using undiscretized training data. At inference time, the policy’s predicted discrete mouse action defines the upper and lower bounds, from which we sample the final mouse action using the fitted truncated normal distribution. Empirically, this approach yields smoother control and better online performance.
@@ 4 Evaluation
[…]
We evaluated our models using (i) the scores from controlled programmatic environments, (ii) human preference evaluations on real games, and (iii) quantitative analyzes of scaling behavior on test loss [^4] Perplexity of the keyboard was used as test loss because some games do not require mouse actions so mouse perplexity was more noisy. and causality. Although several existing game benchmarks exist (Zhang et al., 2025; VIDEO, 2025; Xu et al., 2025; Tomilin et al., 2023), they primarily focus on text-heavy 2D games or sandbox-style 3D environments, which differ substantially from the real-time, first-person gameplay data used in our training setup.
@@
Camera settings can significantly impact performance when playing games in real-time, as overly high sensitivity makes it difficult for the model to adjust to small movements. Detailed camera configurations for DOOM, Quake, and Roblox are provided in Appendix D.1.
@@ 4.1 Simple Programmatic Environment
For Hovercraft, we measure the time (in seconds) required for the agent to complete a full loop. For Simple-FPS, we report the number of hits on the enemy minus the number of hits received. We also report end-to-end inference latency measured on a single RTX 5090 GPU. We compared models of varying sizes trained on the full labeled dataset. Each model was evaluated three times in the same environment, and we report the mean performance. Results are summarized in Table 1. In general, the larger model has worse latency but better capacity.
@@
150M 42 24 80
300M 37 21 64
600M 44 28 62
1.2B 37 37 40
Table 1: Performance on Godot-based programmatic environments across model sizes. The Hovercraft value is the average time (in seconds) that each model takes to finish a loop; the Simple-FPS value is the number of hits of enemy minus the number of hits received by a model. FPS is how many frames per second of inference was feasible on a RTX 5090 GPU.

### v2

@@ Abstract
Behavior cloning has seen a resurgence as scaling model and data sizes demonstrate strong performance. In this work, we introduce an open recipe for training a video game playing foundation model designed for inference in realtime on a consumer GPU. We release all data (8300+ hours of high quality human gameplay), training and inference code, and pretrained checkpoints under an open license. Empirically, we show that our best model achieves performance competitive with human players across a variety of 3D games [^1] All data, code, model checkpoints and videos of game playing are available from the accompanying website https://elefant-ai.github.io/open-p2p/.. We use this recipe to investigate the scaling laws of behavior cloning, with a focus on causal reasoning. In a controlled toy setting, we first demonstrate that increasing training data and network depth leads to the model learning a more causal policy. We then validate these findings at scale, analyzing models up to 1.2 billion parameters. We observe that the causal improvements seen in the toy domain hold true as model size and training steps increase.
@@
[…]
@@ 1 Introduction
In this paper, we train a single model capable of playing a range of 3D video games using raw image observations and producing keyboard and mouse actions in real time on a consumer-grade GPU. Our approach leverages BC with a large-scale dataset collected across a diverse set of games. We release all training and inference code, as well as the dataset (an example is shown in Figure 1), under open licenses. We demonstrate that the model can play simple games that do not require complex planning with a high level of competence.
@@
Causal confusion occurs due to non-causal correlations in the data, which can result in the policy learning to predict the action using these non-causal correlates. An example, taken from De Haan et al. (2019), is that the policy may learn to apply the brakes when it sees the brake lights since these are highly correlated with braking. Obviously, such a non-causal policy performs poorly.
Using both a simple toy example with a rigorous causality metric and empirical results with an approximate causality score, we show that in our setting, scaling both model and dataset size leads to models that more reliably attend to causal signals. These results suggest that a practical approach to mitigating causality issues in behavior cloning is to scale up model capacity alongside data size and diversity.
@@ 2.1 Annotated data
We collect a large-scale, high-quality dataset of human gameplay spanning a diverse set of popular 3D video games. The complete list of games is provided in Appendix A.1. Gameplay is recorded by experienced players who are instructed to capture only active gameplay segments (e.g., excluding lobby or waiting periods). Annotators use a variety of hardware, monitor sizes and resolutions, mouse sensitivities, and gameplay styles, which provide a diverse set of gameplay videos. We do not enforce balanced data distribution across games.
@@
After filtering for quality (see Appendix A.3 for details), the dataset comprises over 8,3008{,}300 hours of high-quality human gameplay, corresponding to around $600$ million image–action pairs. The distribution of recording hours across games is shown in Appendix A.2.
@@ 2.1.1 Text-Annotated Data
Text annotations were generated retrospectively using a commercial VLM. The VLM was prompted to review gameplay segments and infer plausible instructions that could have guided the human player at specific timestamps. This task is inherently challenging, as it requires inferring the player’s underlying intent from recorded behavior and does not admit a unique mapping. While VLMs are not currently capable of playing games at a high level (Zhang et al., 2025), we found that they produce good-quality instructional descriptions when analyzing gameplay videos offline, as shown in Figure 1.
A key challenge in this process is that commercial VLMs typically operate on temporally compressed videos (often around 1 Hz), which is insufficient for text annotation in fast-paced games. To address this limitation, we deliberately prompt the VLM to generate temporally extended, goal-oriented instructions (e.g., “move toward the skull gate”) rather than instantaneous commands (e.g., “turn left”). We further design the prompt to suppress repetitive, high-frequency events, such as shooting, by annotating such behaviors only once per contiguous segment. Finally, the VLM is prompted to generate start and end timestamps, at second-level precision, for each instruction, which are then aligned with the original video recordings.
@@
Figure 2: (2(a)) Architecture of P2P. The core policy transformer and action decoder are both decoder-only transformers. Each timestep begins with a text token $t_{i}$. Since many frames do not contain a text annotation there is a default text token $t_{null}$ used on these frames. This is followed by image token(s) from video frame $o_{i}$ followed by a learnable “reasoning” token $k_{i}$ that grants the model extra computation. The policy transformer then outputs a single action prediction token $a_{in}$. We refer to the reasoning token and the action prediction token as special tokens in the architecture. A smaller transformer, the action decoder, then auto-regressively transforms and samples the single action prediction token into the full action space. Then the true action tokens $a_{i}$ are input so that $a_{in}$ at time $i+1$ can attend to the true action tokens from time ii. (2(b)) Attention mask used in our transformer policy (green denotes 11 and gray 0). This custom mask ensures the action prediction token $a_{in}$ at time $i$ cannot attend to the ground truth action at time ii. Note that no other tokens attend to $a_{in}$ to stabilize the training process.
@@ 2.1.2 Correction Data
To collect correction trajectories, we deploy a trained policy to interact with the game environment while a human annotator monitors its behavior. When the policy encounters out-of-distribution situations (e.g., becoming stuck or exhibiting degenerate behavior), the annotator temporarily takes control to guide the agent back to a valid state. Control is then returned to the policy. For behavior cloning, we used only the human actions and masked out the loss for actions taken by the agent. These correction trajectories are mixed with the original annotated data during training and constitute less than 1% of the total annotated dataset.
@@ 2.1.3 Simple Benchmark Environments
In addition to commercial games, we collected data from two lightweight 3D environments we designed for automated evaluation: Hovercraft and Simple-FPS (see Appendix A.5). These environments are fully programmatic, allowing precise control over game state and difficulty, which provides a fair and controlled benchmark for comparing model performance (see Section 4).
@@ 3 Policy Model
We present a multimodal action policy model, which we refer to as Pixels2Play (P2P). P2P is a text-conditioned policy that takes visual observations and optional textual instructions as input and outputs low-level keyboard and mouse actions. The model builds upon the transformer-based policy architecture introduced in Yue et al. (2025), but introduces substantial extensions to support text conditioning and to improve online performance by explicitly conditioning the backbone transformer on ground-truth action tokens during training.
A primary design constraint is that the model must operate in real time (20 Hz) on high-end consumer GPUs (e.g. NVIDIA RTX 5090), enabling deployment on end-user hardware. To satisfy this requirement, we trained a lightweight, decoder-only transformer backbone from scratch rather than fine-tuning a large pretrained VLM, as explored in prior work (Kim et al., 2024; Tan et al., 2025). This design offers two advantages: (i) it enables a custom image tokenization pipeline that produces a small number of image tokens, allowing the model to attend to longer temporal histories, which is important for games that require long-term memorization; and (ii) the reduced model size and custom architecture ensures fast inference and compatibility with compilation and optimization techniques.
We evaluated several image encoders, including MAGVIT-v2 (Yu et al., 2023; Luo et al., 2024), DINOv2 (Oquab et al., 2023), IBQ (Shi et al., 2025), and CosmosTokenizer (Agarwal et al., 2025). We found that EfficientNet (Tan and Le, 2019), following Pearce and Zhu (2022), provides the best trade-off between representation quality and computational efficiency. Specifically, we use the first six layers of EfficientNet, followed by a linear projection into a small set of visual tokens (typically $N_{i}\in\{1,\dots,4\}$). Unfreezing the image encoder during training consistently improves performance compared to freezing it (Appendix C.1). The model operates directly on the resized raw pixel inputs (192 pixels $\times$ 192 pixels).
Much prior work focused on single-game agents and adopted reduced action spaces that can be modeled as a single categorical variable (Baker et al., 2022; Pearce and Zhu, 2022). In contrast, our setting requires a unified policy capable of operating across many games, necessitating a much larger action space comprising the full keyboard and mouse input domain. We allowed up to four simultaneous key presses and two concurrent mouse actions, making treating the action as a single one-hot combinatorial approach impractical.
@@
For text embeddings, we use EmbeddingGemma (Vera et al., 2025) to encode textual inputs and apply mean pooling over token embeddings to obtain a single text representation. Text embeddings are precomputed, and the text tokenizer is frozen during training.
At each timestep, the policy transformer consumes image tokens ($o_{i}$), a text-conditioning token ($t_{i}$), ground-truth action tokens ($a_{i}$), and a single action prediction token ($a_{\text{in}}$). We further introduce an optional “thinking” token $k$, which provides the model with an additional reasoning step prior to action prediction. We refer to the reasoning token and the action prediction token as special tokens in Figure 2(a). The resulting token count per timestep is $N_{i}+N_{a}+3$. Each token is augmented with a learned type embedding indicating its role (image, text, reasoning, ground-truth action, or action prediction). Rotary positional embeddings (Su et al., 2024) are applied at every transformer layer. During inference, we employ key–value caching with sliding-window attention to bound memory usage.
@@
Behavior cloning often suffers from causal confusion (De Haan et al., 2019). One particular case of this, which becomes worse at higher frequencies, is that the model learns to copy previous actions instead of attending to visual inputs (Wen et al., 2020). Although this issue can be avoided by excluding ground-truth actions from the policy inputs, we find that conditioning the policy on ground-truth action tokens produces more human-like behavior than conditioning on visual observations alone. In particular, the policy learns to sustain actions over multiple frames before switching, more closely resembling human gameplay dynamics. Qualitative differences in behavior are illustrated at https://elefant-ai.github.io/open-p2p/.
@@ 3.1 Mitigating the Training–Inference Gap
Early experiments revealed a substantial performance gap between offline evaluation and online deployment. We traced this gap to discrepancies between training and inference inputs arising from video recording, compression, and resizing. For practical latency and storage considerations, videos undergo two lossy processing steps during data collection: (1) compression and upload on the annotator side, and (2) compressing and resizing to $192\times 192$ for model processing. During inference resizing occurs but no compression takes place.
We observed a substantial divergence in model outputs when using uncompressed raw frames versus resized training frames, which leads to deceptively strong training metrics but poor online performance. We measured this gap by collecting a small number of uncompressed videos. We then compressed the videos using differing compression options and compared the trained model probabilities on the uncompressed versus compressed output. Although compression at the annotation stage is unavoidable (which causes the irreducible gap), we find that the choice of color space during resizing plays a critical role: RGB encoding yields a smaller training–inference gap than YUV encoding (Figure 3(a)). Unfortunately, NVIDIA hardware encoders support only the YUV color space; therefore, we adopt a mixture of QP values ranges from $6$ to $18$ to balance encoding speed and encoding quality.
@@
Data augmentation also substantially reduces the training–inference gap. We applied mild spatial transformations, color perturbations, Planckian jitter, ISO noise, random gaussian or motion blur, sharpening, and translation during training. As shown in Figure 3(b), these augmentations significantly reduced the discrepancy and improved online performance. Consequently, all experiments in this work employ data augmentation. We believe further improvements to reduce the training-inference gap via targeted data augmentation are an interesting area for future work. The detailed parameters are shown in Table 3.
Figure 3: (3(a)) Gap induced by video compression without data augmentation. We measured this gap by comparing model outputs on raw frames (inference) and resized frames (training). An irreducible gap arises from lossy video compression during data collection. The gap was smaller with RGB than YUV encoding and increases as compression quality degrades (lower file size). The x-axis qp denotes the quantization parameter, where larger values indicate lower quality. (3(b)) Data augmentation mitigates this gap when reasonable compression quality is used.
Finally, discretization of the mouse action space can lead to overly aggressive or overly conservative mouse movements. Following Pearce and Zhu (2022), we discretized mouse actions using quantile-based bins, which provide fine resolution near zero but coarse resolution in the tails. To improve robustness, we fit truncated normal distributions to the x and y axes of the mouse action space using undiscretized training data. At inference time, the policy’s predicted discrete mouse action defines the upper and lower bounds, from which we sample the final mouse action using the fitted truncated normal distribution. The mean and standard deviation estimated from the mouse movements data are $\hat{\mu}_{x}=0,\hat{\mu}_{y}=0,\hat{\sigma}_{x}=96,\hat{\sigma}_{y}=22$. Empirically, this approach yields smoother control and better online performance.
@@ 4 Evaluation
We trained policy models at four parameter scales (150M, 300M, 600M, and 1.2B) and compared their performances. More technical details of the training parameters are in Appendix D.1.
We evaluated our models using (i) the scores from controlled programmatic environments, (ii) human preference evaluations on real games, and (iii) quantitative analyses of scaling behavior on test loss22footnotetext: Perplexity of the keyboard was used as test loss because some games do not require mouse actions so mouse perplexity was more noisy. and causality. Although several game benchmarks exist (Zhang et al., 2025; Park et al., 2025; Xu et al., 2025; Tomilin et al., 2023), they primarily focus on text-heavy 2D games or sandbox-style 3D environments, which differ substantially from the real-time, first-person gameplay data used in our training setup.
@@
Camera settings can significantly impact performance when playing games in real time, as overly high sensitivity makes it difficult for the model to adjust to small movements. Detailed camera configurations are provided in Appendix D.2.
@@ 4.1 Simple Programmatic Environment
For Hovercraft, we measure the time (in seconds) required for the agent to complete a full loop. For Simple-FPS, we report the number of hits on the enemy minus the number of hits received. We also report end-to-end inference latency measured on a single RTX 5090 GPU. We compared models of varying sizes trained on the full labeled dataset. Each model was evaluated $16$ times in the same environment, and we report the mean and standard deviation. The results are summarized in Table 1. Overall, the largest model achieves the highest mean score with the lowest variance.
@@
150M $41.1\pm 8.8$ $25.3\pm 6.5$ 80
300M $41.7\pm 8.2$ $25.1\pm 5.6$ 64
600M $38.1\pm 4.2$ $26.5\pm 10.4$ 62
1.2B $36.8\pm 2.9$ $26.6\pm 3.7$ 40
Table 1: Performance on Godot-based programmatic environments across model sizes. For Hovercraft, the score is the time (in seconds) required to complete a loop; for Simple-FPS, the score is the number of enemy hits minus the number of hits received. Each model is evaluated $16$ times per environment, and we report the mean $\pm$ standard deviation. FPS denotes the achievable inference throughput (frames per second) on an RTX 5090 GPU.